Would someone please help me with this I'm just not getting it. Thank you!!
Make sure to supersript all exponents in the final answers should be positive. Final answers should have all single connecting signs and be arranged in descending order.
A) Perform the indicated operations:
(5x^2-3x+2)+(-3x^2+9x-8)-(x^2+3x-10)
B) Use FOIL to multiply:
(3x-8)(x-7)

Answer by stanbon(75887)

A) Perform the indicated operations:
(5x^2-3x+2)+(-3x^2+9x-8)-(x^2+3x-10)
= (5x^2-3x^2-x^2)+(-3x+9x-3x)+(2-8+10)
= x^2 + 3x +4
===============================

B) Use FOIL to multiply:
(3x-8)(x-7)
F: 3x*x = 3x^2
O:3x*-7 = -21x
I:-8*x = -8x
L: -8*-7 = 56
------------------
Combine the terms to get:
= 3x^2 -29x + 56
========================
Cheers,
Stan H.